A textbook store sold a combined total of 348 sociology and math textbooks in a week. The number of math textbooks sold was 64 less than the number of sociology textbooks sold. How many textbooks of each type were sold?
Answer by ikleyn(52817)

.
A textbook store sold a combined total of 348 sociology and math textbooks in a week.
The number of math textbooks sold was 64 less than the number of sociology textbooks sold.
How many textbooks of each type were sold?
~~~~~~~~~~~~~

    Math + Sociology = 348

     x   +   (x+64)  = 348

        2x           = 348 - 64 = 284

         x                      = 284/2 = 142.


ANSWER.  142 Math textbooks and 142+64 = 206 Sociology textbooks.

Solved.